The length of a rectangle is 5 cm more than twice the width. The perimeter of the rectangle is 34 cm. find the dimensions of the rectangle.
Answer by JulietG(1812)

What is a perimeter? It's the area around. Think of a perimeter fence - that's the amount of fence you need to keep predators out of your land. The formula for perimeter is 2W + 2L (2 pieces of fence for each of the lengths and 2 pieces of fence for each of the widths).
.
The problem tells us that the L pieces are (2*W)+5
It also gives us the total, 34. Now we just need to plug in numbers.
.
2W + 2L = 34
2W + 2(2W+5)= 34 (we just substituted the value for L, given in the problem)
2W + 4W + 10 = 34 (distributed the 2 on the outside of the brackets)
Add the Ws:
6W + 10 = 34
Subtract 10 from each side:
6W = 24
Divide each side by 6:
W = 4
.
If the width is 4, we're told that the length is 5 more than twice the width, or (2*4)+5, which is 13.
.
Let's plug the numbers in to make certain.
2W + 2L = 34
2(4) + 2(13) = 34
8 + 26 does indeed equal 34.
Width = 4; length = 13